OpenStudy (anonymous):

how does the parabola x=y*2+3 open? up left right or down?

OpenStudy (anonymous):

up

OpenStudy (anonymous):

no its right @koalamon @bnanglin

OpenStudy (anonymous):

I thought they could not be a function if it did not pass the vertical line test?

OpenStudy (anonymous):

Aperogalics is correct it would be up if the x and y were the other way round in the equation x=y*2+3 is right y=x*2+3 is up

OpenStudy (anonymous):

thanks @DanL33 :)

OpenStudy (anonymous):

he is right, I did not see it was x=y, not y=x

OpenStudy (anonymous):

alright thanks everybody